OpenStudy (anonymous):

Determine whether u and v are orthogonal, parallel, or neither. u=(1,-7) v=(-42,3)

OpenStudy (earthcitizen):

u=1-j7, and v=-42+j3

OpenStudy (anonymous):

Dot product is not zero and both of them are not similar when multiplied by any scalar so they are neither orthogonal nor parallel

OpenStudy (earthcitizen):

\[u*v= (i*i)(j*j)\]

OpenStudy (earthcitizen):

for parallel vectors u*v=|u||v|

OpenStudy (earthcitizen):

|u|=\[\sqrt{a^2+b^2}\]

OpenStudy (earthcitizen):

neither